# Reducing Task Discrepancy of Text Encoders for Zero-Shot Composed Image Retrieval

### v1

@@ 4.1 Experimental setup
Table 2: FashionIQ validation results. The results of RTD combined with Pic2Word [15], SEARLE [16], and LinCIR [17] across different CLIP backbones (ViT-B/32 and ViT-L/14) are shown. Blue denotes the performance gain achieved by our method.
@@
Table 3: CIRR and CIRCO test results. Details are the same as Table 2.
@@
4.2 Main results
Table 2 summarizes the evaluation results on the FashionIQ dataset. In the table, we observe that the incorporation of our approach with ZS-CIR methods significantly improves the performance across all three existing ZS-CIR methods (SEARLE, Pic2Word, and LinCIR) and all backbones (ViT-B/32 and ViT-L/14). For example, regardless of the choice of ZS-CIR methods and backbones, the minimum performance gain for average R@10 and R@50 scores is greater than 2 and 3.5 points, respectively. Table 3 shows a similar trend on the CIRR and CIRCO datasets. Notably, in some metrics on the CIRR and CIRCO datasets, the performance improvements achieved through our method (ViT-B/32) even exceed those obtained by employing a larger backbone (ViT-L/14), which clearly demonstrates the effect of our method. Specifically, in the CIRR R@1 score, SEARLE + RTD (26.29) and LinCIR + RTD (24.82) using ViT-B/32 surpasses the original results of SEARLE (24.89) and LinCIR (23.76) using ViT-L/14. We verify that a similar tendency is observed in the GeneCIS and COCO object composition task datasets, as detailed in the Appendix.
We further evaluate the performance of RTD using the significantly larger backbone (ViT-G/14). In Table 4, by combining RTD and the state-of-the-art ZS-CIR method, LinCIR, we achieve the best ZS-CIR performances on all benchmarks. Details and the full results are provided in the Appendix. We also provide additional qualitative retrieval results in the Appendix.
Table 4: Results on LinCIR [17] with OpenCLIP ViT-G/14 backbone [34].
Method CIRR CIRCO FashionIQ Avg
R@5 R@10 mAP@10 mAP@25 R@10 R@50
LinCIR 64.51 76.12 21.93 24.12 44.53 65.53 49.46
+RTD 67.47 ($\mathbf{+2.96}$) 78.31 ($\mathbf{+2.19}$) 22.29 ($\mathbf{+0.36}$) 24.46 ($\mathbf{+0.34}$) 46.21 ($\mathbf{+1.68}$) 67.26 ($\mathbf{+1.73}$) 50.99 ($\mathbf{+1.53}$)
Table 5: Ablation study. We measure the impact of TCL loss (Eq. 1), refined batch sampling (RB), and refined concatenation scheme (RC) on the validation splits of three CIR datasets. We train RTD based on LinCIR ViT-L/14. The first row equals to the vanilla LinCIR without applying RTD.
TCL RB RC CIRR CIRCO FashionIQ Avg
Text pair Anchor R@5 R@10 mAP@10 mAP@25 R@10 R@50
- - 54.29 67.76 12.67 14.45 27.42 47.71 37.38
$(T_{r},T_{r})$ 55.99 69.72 13.40 15.18 28.16 48.82 38.54
$(T_{r+c},T_{r})$ 58.19 71.54 14.36 16.03 26.93 47.94 39.17
$(T_{r+c},T_{r})$ 58.19 71.27 14.96 16.67 27.42 49.33 39.64
$(T_{r+c},T_{r})$ 54.34 66.97 12.23 13.64 25.02 45.31 36.25
$(T_{r+c},T_{r})$ 57.90 71.13 16.10 17.84 30.24 51.08 40.72
4.3 Ablation studies on the proposed method
Table 5 presents the effectiveness of the proposed components: target-anchored text contrastive loss (TCL), refined batch sampling (RB), and refined concatenation scheme (RC). All evaluation results are on the validation split of CIRR, CIRCO and FashionIQ. All model variants use ViT-L/14 and a projection module $\phi$ from LinCIR [17], making the results in row 1 indicative of the original performance of LinCIR. We first compare the impact of the text pairs fed into TCL loss. We compare our design choice $(T_{r+c},T_{t})$ (from the generated text triplets) with $(T_{r},T_{r})$, which is the sole option for constructing a pair given a single conventional caption $T_{r}$. The results demonstrate that, on average, using generated triplets (3rd row) is more effective than using original conventional text pairs (2nd row), particularly in CIRR and CIRCO. In addition, RB (4th row) and RC (6th row) significantly enhance the overall performance, demonstrating the effectiveness of these components. Finally, we measure the impact of using the frozen text encoder for target caption $T_{t}$, denoted as “Anchor” in the table. Significant performance degradation is observed when the learnable text encoder is used for extracting the embedding of the target caption $T_{t}$ (5th row) compared to the target-anchored case (4th row), supporting the importance of the anchoring design choice.
4.4 More analyses on RTD
In this subsection, we show more analyses on RTD with the same setting to Table 5.
Table 6: Impact of the text triplet generation method. The details are the same as Table 5.
CIRR CIRCO FashionIQ Avg
R@5 R@10 mAP@10 mAP@25 R@10 R@50
Baseline (LinCIR) 54.29 67.76 12.67 14.45 27.42 47.71 37.38
+RTD (LLM-based) 57.90 71.13 16.10 17.84 30.24 51.08 40.72
+RTD (template-based) 56.71 70.34 15.01 16.98 30.37 51.94 40.23
[Impact of the text triplet generation method] There are two variants for the text triplet construction process: LLM-based generation and template-based generation. While our main experiments utilize LLM-based text triplets, we have verified that using rule-based template-based text triplets is sufficient; our template-based text triplets generation process is a fully text-only and automatic process without relying on human labor or LLMs. We describe the details of our template-based generation process in the Appendix A.2. Table 6 shows that the results of our template-based approach are comparable to those obtained with LLM-generated text triplets. We believe this finding implies the simplicity and efficiency of constructing text triplets, underscoring the practicality of RTD.
Table 7: Impact of the update rule. Two update rules are compared: (1) using the original objective from baseline and (2) using RTD. For a fair comparison, in both rules, $\phi$ is updated first and $\psi_{T}$ is updated top on the frozen modules. We use the ViT-L/14 backbone, and numbers are measured on the CIRR dev split, CIRCO validation split, and Fashion IQ validation split.
CIRR CIRCO FashionIQ Avg
R@5 R@10 mAP@10 mAP@25 R@10 R@50
Baseline(Pic2Word) 51.40 64.43 8.77 10.12 25.34 44.88 32.15
+naïve tuning 19.21 27.51 1.29 1.61 4.4 11.15 10.86
+ RTD 56.64 69.77 8.83 9.81 27.59 48.90 36.92
Baseline(LinCIR) 54.29 67.76 12.67 14.45 27.42 47.71 36.86
+naïve tuning 52.67 66.78 11.40 12.99 26.34 45.92 35.52
+ RTD 57.90 71.13 16.10 17.84 30.24 51.08 40.72
@@ 4.4 More analyses on RTD
[Impact of the update rule for the text encoder] To verify that our improvements cannot be achievable solely by tuning the text encoder backbone without considering the task discrepancy, we additionally measure the results of previous methods (Pic2Word and LinCIR) when naively updating text encoders. Namely, after training $\phi$ while keeping all other networks frozen as in previous methods, we additionally update the text encoder using the original loss function, while fixing other modules including $\phi$. We denote this update rule as “naïve tuning” in the Table 7. Unlike RTD, we observe that just naively updating the text encoder (“naïve tuning”) significantly degrades the performance of the baseline. The results indicate that merely updating the text backbone is not beneficial for ZS-CIR; instead, mitigating task discrepancy through RTD is necessary.
@@
[More efficient variants] Table 8 presents the results of the more efficient implementations of our approach in terms of the number of updated parameters. Specifically, instead of updating the entire set of parameters of the text encoder, we explore updating only a few layers of the network when applying RTD, Our findings indicate that updating only the fully connected layers (denoted as “Whole FCs”) nearly matches the performance of the full model while using less than half the number of learnable parameters (40.72 vs. 40.53 average score). Additionally, we verify that updating only three fully connected layers, whose parameter size matches the projection module $\phi$ and constitutes 11.5% of the full model, is also sufficiently effective. We test various three-layer updating strategies: “First 3 FCs”: the first three layers (closest to the input), “middle 3 FCs”: the middle three layers, “Last 3 FCs”: the last three layers, and “Interleave 3 FCs”: an interleaved selection of three layers (first, middle, and last layers). Among these, we verify that the "Interleave 3 FCs" shows the best result, maintaining competitive performance with the full model (40.72 vs. 39.88 average score). We believe these findings suggest a promising direction for enhancing the training efficiency of our approach by selectively updating only specific layers of the text encoder.
5 Discussion and Limitations
As noted by [16, 17], the existing CIR benchmark datasets [1, 18, 19] are somewhat noisy due to the presence of false negatives, leading to unreliable evaluations. A similar issue is reported in the image-text cross-modal retrieval problem by [35, 36]. Specifically, although there is typically one ground truth positive for each query, there may be multiple ground truths within the database. Thus, Gu and Chun et al. [17] primarily use the CIRCO dataset as their main benchmark because CIRCO includes multiple positives and employs a more reliable ranking-based metric, mAP@K [31, 32]. Additionally, they focus on R@K with a larger K (e.g., 10) rather than R@1 in the other benchmarks. In our case, since the overall metrics improve regardless of their type (including mAP@K for CIRCO and R@K with a larger K for the other benchmarks), we believe the improvements from our method are sufficiently reliable despite the noisiness of the evaluation benchmarks.
6 Conclusion
Our research presents RTD, a novel plug-and-play training scheme designed to improve the capabilities of text encoders for ZS-CIR. By leveraging easily obtainable text triplets and implementing target-anchored text contrastive learning, RTD aligns projected and conditioning text embeddings with target embeddings, addressing the challenges posed by task discrepancies in ZS-CIR. Additionally, the integration of hard negatives-based refined batch sampling and a sophisticated concatenation scheme further enhances performance. Empirical evaluations demonstrate that RTD significantly boosts the performance of existing ZS-CIR methods across diverse datasets and model backbones, underscoring its effectiveness and versatility.
Appendix A Additional Implementation Details
A.1 CIR Datasets
We additionally provide results on two more benchmark datasets, GeneCIS [19] and COCO Object Composition introduced by [15], in Section B.1. GeneCIS [19] is also constructed based on COCO images and the Visual Attributes in the Wild dataset [38]. GeneCIS introduces four task variations: (1) focus on an attribute, (2) change an attribute (3) focus on an object and (4) change an object. These tasks explore different aspects of image retrieval and manipulation. For the COCO Object Composition task, we utilize 5000 images from the COCO validation dataset to evaluate object composition. Our objective is to retrieve an image that contains an object specified by a query image, along with scenes or objects described using text. The composed query is constructed by combining “a photo of [\$], [$obj_{1}$], [$obj_{2}$] … and [$obj_{n}$]” where [$obj_{i}$] are text descriptions.
A.2 Training text triplets
In Figure A.1, A.2, examples of both LLM-based and template-based triplets are presented. Both methods generate natural and coherent text.
[Detailed explanation on template-based triplets] Here, we describe the detailed procedure for generating template-based text triplets. We mainly follow the procedure of Compodiff [8], but a text-to-image generation step is not involved. For template-based triplets, we use captions from the CC3M dataset [6] as reference captions $T_{r}$. Firstly, given reference captions, important keywords like nouns are extracted with a part-of-speech (POS) tagger via the Spacy library. Then, the keyword is filtered by frequency filtering with hard thresholding to focus only on frequently occurring keywords. Specifically, we only use keywords that appear more than 100. After applying keyword frequency filtering, the remaining keyword list is used to create caption triplets $(T_{r},T_{c},T_{t})$. To generate the triplets, a keyword from the given $T_{r}$ is selected, and alternative keywords are extracted based on text similarity scores ranging from 0.5 to 0.7, using the SBERT all-MiniLM-L6-v2 [39]. The target caption $T_{t}$ is then constructed by substituting the original keyword with a similar alternative. The conditioning text $T_{c}$ is generated using randomly selected pre-defined templates, as detailed in Table A.1. Here, most of the templates are similar to that of Compodiff [8].
Since the quality of the generated triplets with the above procedure may not be optimal, we employ an additional filtering process. Compodiff [8] employs an additional filtering process that uses cosine similarities between generated images and texts, calculated by CLIP encoders. However, as we do not have images for captions, we filter the inappropriate texts using only textual information inspired by LinCIR [17]. Namely, we calculate the similarity between the CLIP text embedding of $T_{t}$ and the CLIP text embedding of "a photo of [\$]" where [\$] is obtained by $T_{t}$ projected by $\phi$ from LinCIR (ViT-L/14). Following LinCIR noise ($\text{Unif}(0,1)\times\mathcal{N}(0,1)$) is injected before passing through $\phi$. After calculating the above similarity, texts whose similarities are less than the threshold (0.75) are removed. The same process is also applied to the reference caption $T_{r}$ and the intersection of filtering processes for $T_{t}$ and $T_{r}$ is used for the final dataset whose size becomes 1.3M.
Figure A.1: Example of template-based triplet datasets
Figure A.2: Example of LLM-based triplet datasets
Table A.1: The full 50 keyword converting templates
"replace \${source} with \${target}" "substitute \${target} for \${source}"
"apply \${target}" "\${source} is removed and \${target} takes its place"
"convert \${source} to \${target}" "modify \${source} to become \${target}"
"replace \${source} with \${target}" "customize \${source} to become \${target}"
"update \${source} to \${target}" "change \${source} to match \${target}"
"substitute \${target} for \${source}" "\${target} is introduced after \${source} is removed"
"alter \${source} to match \${target}" "\${target} is added in place of \${source}"
"upgrade \${source} to \${target}" "\${target} is introduced as the new option after"
"amend \${source} to fit \${target}" "\${source} is removed and \${target} is added"
"opt for \${target}" "\${source} is removed and \${target} is introduced"
"\${source} is removed" "\${target} is added as a replacement for \${source}"
"add \${target}" "\${target} is the new option available"
"if it is \${target}" "\${target} is added after \${source} is removed"
"\${target} is the updated option" "\${target} is introduced after \${source} is retired"
"\${target} is the updated choice" "tweak \${source} to become \${target}"
"\${source} is replaced with \${target}" "has no \${source}"
"change \${source} to \${target}" "alter \${source} to \${target}"
"swap \${source} for \${target}" "redesign \${source} as \${target}"
"turn \${source} into \${target}" "adapt \${source} to fit \${target}"
"choose \${target} instead of \${source}" "\${target} is the new choice"
"\${target} is the new selection" "exchange \${source} with \${target}"
"transform \${source} into \${target}" "show no \${source}"
"no \${source}" "remove \${source}"
"delete \${source}" "not a \${source}"
"with no \${source}" "without \${source}"
Appendix B Additional experiments
B.1 Results on GeneCIS [19] and COCO object composition
We observe that incorporating our approach with ZS-CIR methods leads to marginal but consistent performance improvements on GeneCIS as shown in Table B.1. The relatively smaller performance difference compared to other datasets can be attributed to the discrepancy between the format of the conditioning text of GeneCIS and the ZS-CIR methods training methodology. Namely, GeneCIS only uses the fixed four text instructions “change attribute”, “focus attribute”, “change object” and “focus object”, which is different from the usual text instruction we expected (e.g., “change the dog to a cat”).
In the experiment on COCO object composition, we observe a significant performance improvement, similar to the results obtained on other datasets in Table B.2. This finding reaffirms that our approach, when combined with ZS-CIR methods, consistently achieves strong performance, demonstrating its generalizability.
Table B.1: GeneCIS results
Average
R@1 R@2 R@3
ViT-B Pic2Word $11.13$ $21.08$ $31.05$
SEARLE $12.19$ $22.56$ $32.03$
+ours $12.82$ ($\mathbf{+0.63}$) $22.97$ ($\mathbf{+0.41}$) $32.44$ ($\mathbf{+0.41}$)
LinCIR $12.23$ $21.29$ $30.80$
+ours $12.83$ ($\mathbf{+0.60}$) $22.83$ ($\mathbf{+1.54}$) $32.22$ ($\mathbf{+1.42}$)
ViT-L Pic2Word $11.18$ $21.45$ $30.55$
+ours $11.92$ ($\mathbf{+0.74}$) $22.32$ ($\mathbf{+0.87}$) $31.33$ ($\mathbf{+0.78}$)
SEARLE $12.30$ $22.08$ $31.29$
+ours $12.40$ ($\mathbf{+0.10}$) $22.82$ ($\mathbf{+0.74}$) $32.37$ ($\mathbf{+1.08}$)
LinCIR $12.45$ $22.66$ $32.06$
+ours $13.18$ ($\mathbf{+0.73}$) $23.12$ ($\mathbf{+0.46}$) $32.77$ ($\mathbf{+0.71}$)
Table B.2: COCO object composition results
COCO
R@1 R@5 R@10
ViT-B Pic2Word $6.88$ $13.6$ $17.52$
+ours $7.62$ ($\mathbf{+0.74}$) $20.23$ ($\mathbf{+6.63}$) $28.69$ ($\mathbf{+11.17}$)
SEARLE $9.52$ $21.45$ $29.38$
+ours $11.01$ ($\mathbf{+1.49}$) $24.34$ ($\mathbf{+2.89}$) $32.84$ ($\mathbf{+3.46}$)
LinCIR $7.15$ $18.38$ $27.3$
+ours $9.59$ ($\mathbf{+2.44}$) $21.66$ ($\mathbf{+3.28}$) $30.66$ ($\mathbf{+3.36}$)
ViT-L Pic2Word $10.26$ $23.67$ $32.53$
+ours $10.26$ ($\mathbf{+0.00}$) $24.66$ ($\mathbf{+0.99}$) $33.56$ ($\mathbf{+1.03}$)
SEARLE $12.07$ $26.13$ $35.17$
+ours $14.38$ ($\mathbf{+2.31}$) $29.74$ ($\mathbf{+3.61}$) $38.09$ ($\mathbf{+2.92}$)
LinCIR $11.37$ $24.53$ $33.85$
+ours $14.6$ ($\mathbf{+3.23}$) $29.84$ ($\mathbf{+5.31}$) $38.87$ ($\mathbf{+5.02}$)
B.2 Results on larger backbone (ViT-G)
We further evaluate the performance of RTD using the significantly larger backbone (OpenCLIP ViT-G/14 [34]). As described in Section 4.3, we use the projection module $\phi$ from LinCIR [17]. Since the pre-trained projection module $\phi$ for LinCIR [17] (ViT-G/14) is not publicly available, we reproduce it and integrate RTD with it. We emphasize that similar to our previous results, RTD again achieves remarkable gains across all datasets. Here, we set the learning rate as $10^{-6}$.
Table B.3: FashionIQ results on larger OpenCLIP ViT-G/14 backbone [34].
Method Shirt Dress Toptee Average
R@10 R@50 R@10 R@50 R@10 R@50 R@10 R@50
LinCIR (reproduced) 46.61 64.72 38.18 60.54 49.26 70.83 44.68 65.36
Table B.4: CIRR and CIRCO results on larger OpenCLIP ViT-G/14 backbone [34].
ViT-G CIRR CIRCO
R@1 R@5 R@10 mAP@5 mAP@10 mAP@25 mAP@50
LinCIR (reported in [17]) $35.25$ $64.72$ $76.05$ $19.81$ $21.01$ $23.03$ $24.18$
LinCIR (reproduced) $34.94$ $64.51$ $76.12$ $20.63$ $21.93$ $24.12$ $25.20$
B.3 Ablations on noise injection
We conduct an ablation study of the different noise types employed for the “refined concatenation scheme” shown in Figure 2. We compare three different noise types, uniform distribution, Gaussian distribution, and LinCIR-ish noise ($\text{Unif}(0,1)\times\mathcal{N}(0,1)$). We also examine the scale of LinCIR-ish noise from 0.1, 0.5, and 1. We report the test scores for CIRR and CIRCO, as well as the FashionIQ validation scores for Pic2Word, SEARLE, and LinCIR in Table D.1 and Table D.2. In the tables, we observe that all noise distributions show decent performance and LinCIR-like noises show slightly better performances than uniform distribution and normal distribution. We also observe that the different scale choice for the LinCIR-like noise somewhat affects the overall performances. In the main experiments, we choose 0.5 for the noise scale, following the observed performance improvements.
Appendix C Qualitative example on CIRCO
We qualitatively illustrate the results of incorporating RTD into LinCIR on the CIRCO dataset in Figure D.1. The visual examples provide an intuitive demonstration of how the integration of RTD enhances the performance of LinCIR, effectively capturing the semantic meaning of the modification descriptions while preserving the relevant visual information from the reference image.
Appendix D Societal Impacts
Figure D.1: Qualitative Results on CIRCO dataset
Table D.1: Noise type variation on CIRR/CIRCO dataset

### v2

@@ B.2 Results on different backbones (ViT-B/32, ViT-L/14)
Table 12: FashionIQ validation results. The results of RTD combined with Pic2Word [36], SEARLE [4], and LinCIR [17] across different CLIP backbones (ViT-B/32 and ViT-L/14) are shown.
@@
Table 13: CIRR and CIRCO test results. Details are the same as Tab. 12.
@@
B.3 Results on larger backbone (ViT-G/14)
As reported in Tab. 2, we evaluate the performance of RTD using the significantly larger backbone (OpenCLIP ViT-G/14 [18]). As described in Sec. 4.3, we use the projection module $\phi$ from LinCIR [17]. Since the pre-trained projection module $\phi$ for LinCIR [17] (ViT-G/14) is not publicly available, we reproduce it and integrate RTD with it. We emphasize that similar to our previous results, RTD again achieves remarkable gains across all datasets. Here, we set the learning rate as $10^{-6}$.
Table 14: FashionIQ results on larger OpenCLIP ViT-G/14 backbone [18].
Method Shirt Dress Toptee Average
R@10 R@50 R@10 R@50 R@10 R@50 R@10 R@50
LinCIR (reproduced) 46.61 64.72 38.18 60.54 49.26 70.83 44.68 65.36
Table 15: CIRR and CIRCO results on larger OpenCLIP ViT-G/14 backbone [18].
ViT-G CIRR CIRCO
R@1 R@5 R@10 mAP@5 mAP@10 mAP@25 mAP@50
LinCIR (reported in [17]) $35.25$ $64.72$ $76.05$ $19.81$ $21.01$ $23.03$ $24.18$
LinCIR (reproduced) $34.94$ $64.51$ $76.12$ $20.63$ $21.93$ $24.12$ $25.20$
@@ B.3 Results on larger backbone (ViT-G/14)
B.4 More efficient variants
@@
Tab. 16 presents the results of the more efficient implementations of our approach in terms of the number of updated parameters. Specifically, instead of updating the entire set of parameters of the text encoder, we explore updating only a few layers of the network when applying RTD, Our findings indicate that updating only the fully connected layers (denoted as “Whole FCs”) nearly matches the performance of the full model while using less than half the number of learnable parameters (40.72 vs. 40.53 average scores). Additionally, we verify that updating only three fully connected layers, whose parameter size matches the projection module $\phi$ and constitutes 11.5% of the full model, is also sufficiently effective. We test various three-layer updating strategies: “First 3 FCs”: the first three layers (closest to the input), “middle 3 FCs”: the middle three layers, “Last 3 FCs”: the last three layers, and “Interleave 3 FCs”: an interleaved selection of three layers (first, middle, and last layers). Among these, we verify that the “Interleave 3 FCs” shows the best result, maintaining competitive performance with the full model (40.72 vs. 39.88 average scores). We believe these findings suggest a promising direction for enhancing the training efficiency of our approach by selectively updating only specific layers of the text encoder.
B.5 Effectiveness of RTD across dataset scales
We conduct experiments with various scales of training text triplets. For the small-scale text triplets, we sub-sampled text triplets from LLM-based text triplets. Thus, the last row in the Tab. 17 denotes the original result (LLM-based RTD result). We also measure the effectiveness of RTD using large-scale text triplets (up to 5M) by combining publicly available text triplets (IP2P, CoVR) with ours (LLM-based, rule-based). Here, validation splits of all three benchmark datasets are utilized and full results will be included in the final version.
As shown in Tabs. 17 and 18, we confirm that small-scale text triplets are sufficient to achieve the effectiveness of RTD. We believe the main reason for this is that, to reduce task discrepancy, only the relationship between the concatenated caption (reference caption + conditioning caption, $T_{r+c}$) and the target caption $T_{t}$ needs to be learned. We believe this learning task requires much less data compared to learning representations from scratch. Moreover, since the text encoder is already pre-trained, the model does not need significant changes to learn this simple but crucial learning task for CIR.
Table 17: Results across different scales of LLM-based text triplets. In each row, text triplets are sub-sampled from 2.5M original LLM-based text triplets provided by Compodiff [16]
# of triplets CIRR R@5 CIRCO mAP@10 FashionIQ R@10 Avg
1K 56.64 15.66 29.89 34.06
50K 57.40 15.95 30.77 34.71
100K 57.16 16.03 30.57 34.59
2.5M 57.90 16.10 30.24 34.75
Table 18: Results of larger-sized text triplets
IP2P CoVR Compodiff Template-based CIRR R@5 CIRCO mAP@10 FashionIQ R@10 Avg # of triplets
58.65 15.94 29.62 34.74 450k
59.82 15.35 29.58 34.92 700k
57.90 16.10 30.24 34.75 2.5M
56.71 15.01 30.37 34.03 1.3M
59.32 16.10 30.81 35.41 1.25M
59.08 16.15 30.97 35.40 3.75M
58.65 16.54 31.22 35.47 5.05M
B.6 Ablations on filtering process
In rule-based text triplet generation, we highlight that the filtering process using the projection module from LinCIR is marginally effective. As demonstrated in the Tab. 19, even without the filtering procedure, the enhancement of RTD from LinCIR remains considerable. This result demonstrates that the effectiveness of our rule-based text triplets is not solely dependent on the use of the projection module from LinCIR in the filtering process.
Table 19: Ablations on filtering process
Type LinCIR-based filtering CIRR R@5 CIRCO mAP@10 FashionIQ R@10 Avg
LinCIR - 54.29 12.67 27.42 31.46
+RTD (rule-based) 55.49 14.75 30.24 33.49
+RTD (rule-based) 56.71 15.01 30.37 34.03
Table 20: Noise type variation on CIRR/CIRCO dataset